OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is −6, the last term is −24,576, and the common ratio is 4?

OpenStudy (texaschic101):

sn = a1(1 - r^n) / (1 - r) n = the sum of n terms = 7 a1 = first term = -6 r = common ratio = 4 now we sub ... s7 = -6(1 - 4^7) / (1 - 4) s7 = -6(1 - 16384) / -3 s7 = -6( - 16383) / -3 s7 = 98298/-3 s7 = - 32766

OpenStudy (texaschic101):

this should help... http://www.regentsprep.org/Regents/math/algtrig/ATP2/GeoSeq.htm